### v4

# Low Variance Off-policy Evaluation with State-based Importance Sampling

Abstract
In off-policy reinforcement learning, a behaviour policy performs exploratory interactions with the environment to obtain state-action-reward samples which are then used to learn a target policy that optimises the expected return. This leads to a problem of off-policy evaluation, where one needs to evaluate the target policy from samples collected by the often unrelated behaviour policy. Importance sampling is a traditional statistical technique that is often applied to off-policy evaluation. While importance sampling estimators are unbiased, their variance increases exponentially with the horizon of the decision process due to computing the importance weight as a product of action probability ratios, yielding estimates with low accuracy for domains involving long-term planning. This paper proposes state-based importance sampling, which drops the action probability ratios of sub-trajectories with “negligible states” – roughly speaking, those for which the chosen actions have no impact on the return estimate – from the computation of the importance weight. Theoretical results show this reduces the ordinary importance sampling variance from $O(\exp(H))$ to $O(\exp(X))$ where $X<H$ is the largest subtrajectory with non-negligible states. To identify negligible states, two search algorithms are proposed, one based on covariance testing and one based on state-action values. We formulate state-based variants of ordinary importance sampling, weighted importance sampling, per-decision importance sampling, incremental importance sampling, doubly robust off-policy evaluation, and stationary density ratio estimation. Experiments in four distinct domains show that state-based methods consistently yield reduced variance and improved accuracy compared to their traditional counterparts.
1 Introduction
In reinforcement learning (RL) [18], an agent explores an environment which provides reward signals denoting the desirability of the behaviour, and through this feedback, it learns a policy to optimise a utility function, such as the expected cumulative reward, in that environment. In many domains, including medical interventions (e.g. surgery, diabetes or sepsis), financial decision-making (e.g. advertising or trading), and safe navigation, the exploration process will be too costly as it requires trying out suboptimal policies. In such domains, ideally one would be able to use expert decision-making as an initial starting point and improve the decision-making policy based on a collected data set rather than online interaction with trial-and-error. Such domains motivate the use of off-policy evaluation [25, 7], a class of methods which estimate the utility of the policy to be evaluated using trajectories sampled from a behaviour policy that can be safely used for exploration.
Importance sampling (IS) is a traditional statistical technique which estimates an integral over one distribution of interest while only having samples from another distribution. The expected cumulative reward, discounted or undiscounted, can be seen as such an integral, and therefore IS is an important technique for off-policy evaluation [15, 22, 2]. Such techniques have been used successfully in the context of diabetes treatment [2] and digital marketing [22]. Unfortunately, due to computing the importance weight as a product of action probability ratios across the entire episode, the variance of IS based off-policy evaluation grows exponentially with the horizon $H$ of the decision process, a problem also known as “the curse of horizon” [8]. This makes IS based off-policy evaluation perform more poorly in domains requiring long-term planning (e.g. safe navigation).
To reduce the variance of the importance sampling estimator, Guo et al. (2017) [5] introduce covariance testing, which drops terms (i.e. action probability ratios) from the importance weights based on the expected value of the product of the terms dropped being equal to 1 and the covariance between the estimate and the dropped product being equal to 0. This leads to the implementation of incremental importance sampling (INCRIS), which drops all but the most recent action probability ratios from the importance weight of the per-decision importance sampling (PDIS) estimator [15].
This paper takes a different approach to reduce the variance of importance sampling estimators, which we call state-based importance sampling (SIS). While INCRIS drops all but the $k$ most recent action probability ratios, there is not always an a priori reason why the number $k$ should be related across trajectories. That is, the $i$’th decision in one trajectory is often unrelated to the $i$’th decision in another trajectory – e.g. due to stochasticity of the starting state, the policy, or the transition dynamics – and therefore the $k$ previous time steps may not be the best criterion for dropping action probability ratios. Instead, by dropping action probability ratios associated with a particular state or set of states, the dropped terms for SIS have a related meaning across trajectories. More specifically, SIS drops the action probability ratios corresponding to states that do not affect the estimate when dropped any time they occur in a trajectory; we call such states “negligible states” (see Section 4). We formulate algorithms for identifying negligible states automatically based on covariance testing as well as state-action values (see Section 5). The state-action value based formulation is especially convenient to integrate with other variance reduction techniques, so we demonstrate state-based approaches to a variety of low-variance importance sampling estimators in addition to traditional IS. Beyond ordinary importance sampling, there are also state-based formulation of other importance sampling estimators, doubly robust estimation, and stationary density ratio estimation (see Section 6). Section 7 details the experiments conducted in a deterministic and stochastic lift domain, which include states for which actions do not affect the next state nor the next reward, as well as inventory management and taxi domains as more generic RL benchmarks. Results demonstrate that state-based approaches consistently yield improved accuracy when compared to their traditional counterparts.
2 Problem statement and preliminaries
This present paper is based on finite-horizon undiscounted MDPs, which can be defined based on the tupe $(\mathcal{S},\mathcal{A},r,\mathcal{T})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is a reward function, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\to\Delta^{\mathcal{S}}$ is the transition dynamics model defined over the probability simplex over the state space, $\Delta^{\mathcal{S}}$. The goal of the reinforcement learning agent is to select actions that maximise the expected cumulative reward, or “return” for short. The decision-making is represented by the policy of the agent, which takes the form $\pi:\mathcal{S}\to\Delta^{\mathcal{A}}$; that is, the policy $\pi$ is a function outputting a vector of action probabilities for a given state.
Off-policy evaluation computes the expected return $\mathcal{G}$ of an evaluation policy $\pi_{e}$ based on samples from a behaviour policy $\pi_{b}$. This can be done using importance sampling,
$\mathcal{G}$ $=\mathbb{E}_{\pi_{e},\mathcal{T}}\left[\sum_{t=1}^{H}r_{t}\right]$
$=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[\sum_{t=1}^{H}r_{t}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)\right]\,,$
which samples from $\pi_{b}$ and then corrects for the probability under $\pi_{e}$.
A long-standing problem with importance sampling in this manner is that if the horizon is large, then the variance on the estimate of $\mathscr{G}$ is prohibitively large. To see why, define the estimate $\hat{G}$
$\hat{G}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}^{(i)}|s_{t}^{(i)})}{\pi_{b}(a_{t}^{(i)}|s_{t}^{(i)})}\right)$
where $i$ indexes the trajectory of the state $s_{t}^{(i)}$, action $a_{t}^{(i)}$, and reward $r_{t}^{(i)}$ at time $t$. Noting that $\hat{G}$ is the sample mean of the random variable $X=\sum_{t=1}^{H}r_{t}\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)$ and assuming $r\in[0,r_{\text{max}}]$, the following provides a general upper bound for the variance of this estimator:
$\text{Var}(\hat{G})$ $=\frac{1}{n}\text{Var}(X)$
$\leq\frac{1}{4n}\left(\sum_{t=1}^{H}r_{\text{max}}\left(\prod_{t=1}^{H}\rho_{\text{max}}\right)\right)^{2}$
$=\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{H}\right)^{2}\,,$
where $\rho_{\text{max}}=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\pi_{e}(a|s)/\pi_{b}(a|s)$ and the second line follows from assuming rewards in $[0,r_{\text{max}}]$ and applying Popoviciu’s inequality [13], which states that $\text{Var}(X)\leq(M-m)^{2}/4$ for any random variable $X$ with range $[m,M]$. In episodic environments, one may also replace $H$ with the maximal episode length as the effective horizon. In addition to the general upper bound, it is also possible show an exponential lower bound under certain assumptions. For instance, under the assumptions that the Markov chain $\{s_{t},a_{t}\}$ is Harris recurrent, has the geometric drift property, and that $|\log(\rho_{t}|)|<\infty$ for $\rho_{t}=\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}^{(i)}|s_{t})}{\pi_{b}(a_{t}|s_{t})}$, Liu et al. (2020) [9] show that the variance is lower bounded by
$\text{Var}(\hat{G})\geq\frac{G^{2}}{4}\exp\left(\frac{Hc^{2}}{8c_{1}^{2}||B||_{\infty}}\right)-G^{2}\,,$
where $c=\mathbb{E}_{d^{\pi_{b}}}\left[KL(\pi_{b}(a|s)||\pi_{e}(a|s))\right]$ based on the average state-visitation distribution $d^{\pi_{b}}$; $c_{1}$ is a positive constant; and $B$ is the test function of the geometric drift.
3 Related work
In addition to the earlier-mentioned INCRIS algorithm, SIS is related to three lines of work for low-variance importance sampling, namely stationary density ratio estimation, marginalised importance sampling, and conditional importance sampling. The section below briefly discusses these methods followed by a brief comparison to SIS. The section then concludes by referring to other approaches for off-policy evaluation and briefly mentioning important aspects about the application context of off-policy evaluation.
3.1 Approaches related to state-based importance sampling
Stationary density ratio estimation (SDRE) [8] avoids the cumulative product across time and instead is based on the average visitation distribution of state-action pairs. Denoting the average visitation distribution of a state $s_{t}^{(i)}$ as $d_{t}^{\pi}(s_{t}^{(i)})$ and its estimate as $\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})$, SDRE estimates the return as
$\hat{G}_{SDRE}=\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\frac{\hat{d}^{\pi_{e}}(s_{t}^{(i)})}{\hat{d}^{\pi_{b}}(s_{t}^{(i)})}\,,$ (1)
thereby removing the curse of horizon by transforming an exponential dependence on $H$ into a polynomial dependence on $H$. The approach is thereby suitable for large infinite-horizon problems. Unfortunately, the approach requires a suitable model (e.g. a neural network) for estimating the true stationary density ratio. While the approach designs a mini-max loss function to train the model, the accuracy of the approach in estimating the return depends on the accuracy of the model (at worst, the MAE may be equal to the loss). The approach requires assumptions of ergodicity and the state space being finite, and due to requiring a stationary distribution, the approach may not work on finite-horizon MDPs.
Marginalised importance sampling (MIS) [26] is closely related to SDRE. The MIS estimator is given by
$\hat{G}_{MIS}=\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{H}\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})\frac{\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})}{\hat{d}_{t}^{\pi_{b}}(s_{t}^{(i)})}\,,$ (2)
where one can note the difference with SDRE by indexing the state distribution $\hat{d}_{t}^{\pi_{e}}$ at a particular time $t$ as well as estimating the reward $\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})$ for a particular time $t$. The estimator of the state distribution is updated according to $\hat{d}_{t}^{\pi_{e}}=\hat{\mathcal{T}}_{t}^{\pi_{e}}(s_{t}|s_{t-1})\hat{d}_{t-1}^{\pi_{e}}$, where $\hat{\mathcal{T}}_{t}^{\pi_{e}}(s_{t}|s_{t-1})$ is the empirical importance-weighted visitation of $s_{t-1}\to s_{t}$. The estimator $\hat{r}_{t}^{\pi_{e}}$ is the importance-weighted reward for transitions $s_{t-1}\to s_{t}$. Assumptions are 1) bounded expected value and variance of reward and variance; 2) the behaviour policy has support for all $t$ and all $s_{t}$; and 3) the density ratios and importance weights are bounded. The approach yields a variance that is polynomial in $H$, according to $O(H^{3})$ in general. Looking at specific cases, one can reduce to $O(H^{2})$ when $\pi_{e}\approx\pi_{b}$, and $O(H)$ when the maximal value does not depend on the horizon. Within the formalism of a tabular MDP, i.e. where state space and actions space are limited such that each state-action pair can be frequently visited, Tabular MIS [28] provides an alternative approximation to $\hat{r}_{t}^{\pi_{e}}(s_{t}^{(i)})$ and $\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})$ based on the observed visitation frequencies. Tabular MIS provides a lower bound in line with the Cramer-Rao bound of $\Omega(H^{2}/n)$.
Conditional importance sampling (CIS) uses conditional Monte Carlo over the importance weight to yield lowered variance. Denoting $\rho_{1:t^{\prime}}=\prod_{t=1}^{t^{\prime}}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}$ and $\phi_{t}$ as a statistic obtained from the trajectory, CIS can be written as
$\hat{G}_{CIS}=\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\mathbb{E}[\rho_{1:H}|\phi_{t}]]\,,$ (3)
which yields $\text{Var}(G\mathbb{E}[\rho_{1:H}|\phi_{t}])\leq\text{Var}(G\rho_{1:H})$. Liu et al. (2020) [9] make use of this property to better understand IS, PDIS, and marginalised IS, noting that in this framework the PDIS estimate has $\phi_{t}=s_{1:t},a_{1:t}$ and the marginalised IS estimate has $\phi_{t}=s_{t},a_{t}$. Their results show several findings:
  • For finite horizon, if $\rho_{1:t}$ and $\rho_{1:t}r_{1:t}$ are positively correlated, then PDIS’ variance is lower than traditional importance sampling. The positive correlation happens when the target policy is near-optimal (because then high-density trajectories will be associated with large rewards).
  • For finite horizon, for $t\leq k$, $\text{Cov}(\rho_{1:t}r_{t},\rho_{1:t}r_{k},)\leq\text{Cov}\left(\frac{\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})}{\hat{d}_{t}^{\pi_{b}}(s_{t}^{(i)})}r_{t},\frac{\hat{d}_{t}^{\pi_{e}}(s_{t}^{(i)})}{\hat{d}_{t}^{\pi_{b}}(s_{t}^{(i)})}r_{k}\right)$, then $\text{Var}(\hat{G}_{PDIS})\leq\text{Var}(\hat{G}_{MIS})$.
  • For infinite horizon, $H\to\infty$, and under some conditions related to ergodicity and recurrence, PDIS has variance $O(H^{2})$ and MIS with oracle importance weights has $O(H^{2})$ but in practice requires approximation of the importance weights, leading to $O(H^{2}(1+\epsilon))$. However, in the general case, these are exponential rather than polynomial.
Rowland et al. (2019) [17] present a related framework for conditional importance sampling with emphasis on trajectories given particular start state-action pair. They introduce novel algorithms arising within the framework, namely return-conditioned, reward-conditioned, and state-conditioned importance sampling. When the equations cannot be solved analytically, the technique often comes with a requirement to form an importance weight regression model, which may increase variance and reduce accuracy.
Compared to the above, SIS estimators can be integrated with a variety of techniques, including ordinary IS, INCRIS, PDIS, and SDRE, which means that it inherits the assumptions of the base-estimator (e.g. with or without function approximation, with stationary distribution assumptions or not). The key property of SIS based estimators is that they work by removing states that are negligible (see Definition 1) from the computation of the importance weight. As such they require a suitable negligible state set as input which is based on either a search algorithm using approximations or prior knowledge. SIS is expected to perform particularly well when one has many states which do not contribute to the expected return and when matched with a base-estimator that is suitable for the domain. The upper bounds for the variance and the MSE for SIS provided in equations 5 and 6 illustrate the effect it can have in case ordinary IS is used as a base-estimator. The integration with other base-estimators is discussed in Section 6, and Section 7 reports experiments comparing state-based variants to their traditional counterparts.
3.2 Other approaches to variance reduction in off-policy evaluation
Other techniques have proven to be useful for off-policy evaluation, and their integration with state-based importance sampling is demonstrated in Section 7. Weighted importance sampling (WIS) and per-decision importance sampling (PDIS) [15] reduce the variance through weighting by normalising based on the sum of trajectory weights and by only accounting for time steps before the decision at the time the action was taken, respectively. The direct method models the reward function which is conceptually simple but requires a model, the accuracy of which often depends on the evaluation policy [3], and any statistical bias in the reward model will be reflected in the estimator. Doubly robust methods [3, 6, 4, 21] use estimators that combine the direct method and importance sampling based on traditional statistical techniques for missing data (e.g. [1]). The doubly robust estimator benefits from the low variance of the direct method and the low bias – or unbiasedness, if the behaviour policy is known – of importance sampling methods. Guo et al. [5] combine INCRIS with options [19], temporally extended actions that help to effectively reduce the horizon of the importance sampling estimator. Rather than competing with these methodologies, state-based methods for importance sampling can be seen as a supplement to these; using the Q-value based identification method (see Section 5), it is straightforward to implement this integration and Section 7 demonstrates this typically reduces the MSE of these approaches.
Other variance techniques not investigated in this paper are also of further interest. Similar to INCRIS, we expect that SIS can also be combined with options for improved accuracy in long-horizon problems. Multiple importance sampling [12] is a traditional Monte Carlo technique that has also been investigated in the context of off-policy evaluation [11]. Estimator selection [25] shares some similarities with state-based importance sampling in the fact that it also selects an estimator by comparing the MSE estimates for different estimators.
3.3 Applying off-policy evaluation
Off-policy evaluation has two primary uses. First, when few trajectories are available and the designer wants to make a decision on whether or not to use the evaluation policy instead of a given expert policy; primary examples include medical treatment [2] and digital marketing [20], which would incur significant costs and safety risks when the evaluation policy’s performance is not estimated correctly. Second, off-policy evaluation can be used to optimise the policy in what is called off-policy improvement. Important topics in this area are how to ensure monotonic improvement and how to provide confidence estimates on improvement. Off-policy improvement is based on batch algorithms since these take into account the trajectories sampled from the behaviour policy. High-confidence policy improvement (HCOPE) [24] uses confidence estimates on the evaluation [23] to ensure with high probability that the new policy improves on the old. Policy optimisation via importance sampling (POIS) [10] optimises surrogate objective that effectively captures the trade-off between the estimated improvement and the uncertainty due to importance sampling using the 2-Renyi divergence of the evaluation policy and the behaviour policy. POIS benefits from using multiple importance sampling and per-decision importance sampling [11] so a possible integration with state-based methods may also be considered in future work. Uniform OPE [27] provides stronger guarantees for off-policy evaluation by simultaneously evaluating all policies in a policy class, thereby integrating naturally with policy improvement.
4 State-based Importance Sampling
State-based importance sampling (SIS) mitigates the above issue by constructing an estimator $\hat{G}_{SIS}$ that selectively drops the action probability ratios of selected states from the product to compute the importance weight, such that 1) on expectation, the dropped term is equal to 1; and 2) the dropped term does not covary with the estimator $\hat{G}_{SIS}$. Additionally, we note that the variance upper bound is lowered if the remaining action probability ratios have a low upper bound, although this does not affect the estimator’s bias. The theory behind this is outlined below.
Let $\mathcal{S}^{A}\subset\mathcal{S}$ and $\mathcal{S}^{B}=\mathcal{S}\setminus\mathcal{S}^{A}$ its complement. Let $G=\sum_{t=1}^{H}r_{t}$. Also for any given trajectory $\tau=\{s_{1},a_{1},\dots,s_{H},a_{H}\}$ taken by a policy $\pi_{b}$ on an MDP $\mathcal{M}$, let
$A=\begin{cases}\prod_{t\in[H]:s_{t}\in\mathcal{S}^{A}}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})},&\text{if $\exists s\in\mathcal{S}^{A}\cap\tau$}\\ 1,&\text{otherwise}\\ \end{cases}$
and let
$B=\begin{cases}\prod_{t\in[H]:s_{t}\in\mathcal{S}^{B}}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})},&\text{if $\exists s\in\mathcal{S}^{B}\cap\tau$}\\ 1,&\text{otherwise,}\\ \end{cases}$
where $[H]=\{1,2,\dots,H\}$. Then, following the covariance testing approach by Guo et al. (2017) [5], the expected return can be rewritten as
$\mathscr{G}$ $=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[G\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)\right]$
$=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[GAB\right]$
$=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[A\right]\mathbb{E}_{\pi_{b},\mathcal{T}}\left[BG\right]+\text{Cov}(A,BG)\,,$
where the last step follows from $\text{Cov}(X,Y)=\mathbb{E}[XY]-\mathbb{E}[X]\mathbb{E}[Y]$. The above implies that when provided with a sufficient number of samples from the random variable $BG$, one can accurately estimate $\mathscr{G}$ based on samples of $BG$ whenever $\mathbb{E}_{\pi_{b}}\left[A\right]=1$ and $\text{Cov}(A,BG)=0$.
Condition 1: $\mathbb{E}_{\pi_{b}}\left[A\right]=1$ This condition holds if
$\mathbb{E}_{\pi_{b},\mathcal{T}}\left[\prod_{t\in[H]:s_{t}\in\mathcal{S}^{A}}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right]=1\,.$
The transition dynamics model $\mathcal{T}$ of the MDP is not known and therefore the frequency of states in $\mathcal{S}^{A}$ occurring in these trajectories is unknown and may vary considerably. However, the random variable $A$ can be seen as the importance weight of an evaluation policy $\pi_{e}$ that is equal to $\pi_{e}$ for $s\in\mathcal{S}^{A}$ and equal to $\pi_{b}$ for $s\in\mathcal{S}^{B}$. Therefore, $A$ can be seen as an importance weight of the entire trajectory under a different evaluation policy $\pi_{e}^{A}$, which we will call the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$. Proof is given in Appendix A.
Condition 2: $\text{Cov}(A,BG)=0$ This condition holds if $BG$ does not grow bigger (resp. smaller) if $A$ increases (resp. decreases). Condition 2 can be verified according to the algorithm proposed in Section 5; while the algorithm verifies it explicitly, it may also be done implicitly due to the contribution to the MSE.
Condition 3: the states $s\in\mathcal{S}^{B}$ have low maximal action probability ratio.
Provided that for some $\epsilon>0$, $\pi_{b}(s,a)\geq\epsilon$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ and that $\pi_{e}\leq 1$, a general upper bound to the action probability ratio is $\rho_{\text{max}}=1/\epsilon^{H}$. Therefore, this condition is satisfied in general when a uniform random policy is chosen for the behaviour policy $\pi_{b}$, since these have $\epsilon=1/|\mathcal{A}|$. This may not always desirable, e.g. when requiring a safe expert policy, but for variance reduction in SIS this can be an important factor for improving worst-case bounds. Note that Condition 3 has a beneficial effect on variance reduction but has no influence on the bias. A minimalistic interpretation of Condition 3 would to have a soft behaviour policy (i.e. with no zero-probability actions) such that no importance ratios can be infinite, which is a typical assumption in importance sampling approaches to off-policy evaluation.
Under the above conditions, the SIS estimator is constructed for a given dropped state set $\mathcal{S}^{A}$ as
$\hat{G}_{SIS}(\mathcal{S}^{A})=\frac{1}{n}\sum_{i=1}^{n}G^{(i)}\left(\prod_{t\in[H]:s_{t}^{(i)}\in\mathcal{S}^{B}}\frac{\pi_{e}(a_{t}^{(i)}|s_{t}^{(i)})}{\pi_{b}(a_{t}^{(i)}|s_{t}^{(i)})}\right)\,.$ (4)
Rather than exponential dependence on $H$, this variance upper bound of $\hat{G}_{SIS}$ is exponentially dependent on the maximal number of occurrences of states in $\mathcal{S}^{B}$. Denoting the subtrajectory $\tau_{B}:=\{t\in[H]:s_{t}^{(i)}\in\mathcal{S}^{B}\}$ and $M_{B}:=\max_{\tau\in\tau_{B}}|\tau|$, we have
$\text{Var}(\hat{G}_{SIS}(\mathcal{S}^{A}))\leq\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}\,.$ (5)
This result implies that the performance of the SIS estimator will be significantly improved if $M_{B}$ is sufficiently small, i.e. if state dropping is a frequent occurrence.
The underlying principle behind covariance testing [5] is that dropping a term $A$ from the importance-weighted return $ABG$ for an estimate $BG$ introduces a bias that is equivalent to the covariance $\text{Cov}(A,BG)$. Consequently, the MSE is given by
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))=\text{Var}(\hat{G}_{SIS}(\mathcal{S}^{A}))+\text{Cov}(A,BG)^{2}\,.$ (6)
This leads to the definition of negligibility below.
Definition 1.
$(\pi_{e},\pi_{b},\epsilon)$-decomposable MDP and $\epsilon$-negligible state set. Let $\epsilon>0$, $\mathcal{M}$ be an MDP $\langle\mathcal{S},\mathcal{A},r,\mathcal{T}\rangle$, and let $\pi_{e}$ and $\pi_{b}$ two policies in $M$. Then $\mathcal{M}$ is $(\pi_{e},\pi_{b},\epsilon)$-decomposable iff there exists a set $\mathcal{S}^{A}$ such that $\text{Cov}(A,BG)<\epsilon$. Any such set $\mathcal{S}^{A}$ is called an $\epsilon$-negligible state set for the off-policy evaluation problem $\langle\mathcal{M},\pi_{e},\pi_{b}\rangle$.
When $\mathcal{S}^{A}$ is an $\epsilon$-negligible state set for an off-policy evaluation problem, this implies via Equation 56 that
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))\leq\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}+\epsilon^{2}\,.$ (7)
5 Optimising the dropped state set
We propose two search algorithms to optimise the dropped state set, one based on covariance testing and one based on state-action values (a.k.a. Q-values). Each search algorithm comes with its own identification of negligible states, i.e. those that are believed to be allowed to drop with only a negligible increase in bias. Inaccurate identification of negligible state sets can increase the MSE significantly. However, as illustrated in Appendix B, if many states can be dropped then despite this additional bias, the performance of SIS is often a significant improvement over IS.
5.1 Covariance testing
To identify $\epsilon$-negligible state sets, one can take the approach taken in incremental importance sampling (INCRIS) [5], which is to compute an empirical estimate of the mean squared error (MSE) based on the bias-variance decomposition of the MSE:
$\widehat{\text{MSE}}(\hat{G}_{SIS}(\mathcal{S}^{A}))=\widehat{\text{Var}}(\hat{G}_{SIS}(\mathcal{S}^{A}))+\widehat{\text{Cov}}(A,BG)^{2}\,,$ (8)
where $\widehat{\text{Cov}}(A,BG)$ is the empirical covariance estimate and $\widehat{\text{Var}}(A)$ is the empirical variance estimate. In SIS, the state set $\mathcal{S}^{A}$ to be dropped from the importance weight computation may require checking many subsets of states within $\mathcal{S}$. In this paper, we perform experiments in a limited state space which allows performing an exhaustive search over many subsets – although rather than the full power set, we limit the search to sets with cardinalities of at most 2. The algorithm defines the best state set incrementally by defining a state set $\texttt{best}\leftarrow\mathcal{S}^{A}$ whenever either $\widehat{\text{MSE}}(\mathcal{S}^{A})<\widehat{\text{MSE}}(\texttt{best})$ or $\widehat{\text{MSE}}(\mathcal{S}^{A})<\widehat{\text{MSE}}(\texttt{best})*(1+\epsilon_{s})\land|\mathcal{S}^{A}|>|\texttt{best}|$ based on a user-defined $\epsilon_{s}>0$. Before doing so the state set must satisfy Condition 2, which is verified by checking $\epsilon$-negligibility using the estimated covariance by $\widehat{\text{Cov}}(A,BG)<\epsilon$.
5.2 Q-value based identification
Another option we explore as a search algorithm to identify negligible states is to keep track of a state-action value table which indicates the expected cumulative reward from a particular state-action pair. The reasoning is that, if the state’s action probability ratio is to be dropped, it should have state-action values that are highly similar. To do this end, a model $M$ of the MDP is formulated based on the same trajectories that are given for off-policy evaluation. The model estimates, based on the observed state visitation frequencies and reward intake, for each $s\in\mathcal{S}$ the initial distribution $\hat{P}_{1}(s)$ and for each $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ the transition probability $\hat{\mathcal{T}}(s,a,s^{\prime})$ as well as the estimated average reward $\hat{r}(s,a)$. Using the model, one can then compute time-specific Q-values via stochastic dynamic programming:
$\hat{Q}_{t}(s,a)=\hat{r}(s,a)+\sum_{s^{\prime}\in\mathcal{S}}\hat{\mathcal{T}}(s,a,s^{\prime})\pi_{e}(:|s^{\prime})\cdot\hat{Q}_{t+1}(s^{\prime},:)\,,$ (9)
where :::: indicates vectorisation.
Intuitively, the action probability ratios for a state for which actions have similar Q-values can be ignored for the purposes of computing $\hat{G}$ since the same cumulative value would be obtained regardless of the action. This leads to the Q-value based algorithm, which identifies a state $s\in\mathcal{S}$ as negligible if for all $t\in[H]$, $|\hat{Q}_{t}(s,a)-\hat{Q}_{t}(s,a^{\prime})|<\epsilon$ for all $a,a^{\prime}\in\mathcal{A}$.
More formally, a simple proof can show that using this criterion leads to an SIS estimator with limited absolute bias depending on the user-defined $\epsilon$, the maximal number of dropped weights, and the model approximation error $\epsilon_{2}$ (see Appendix C). Consistent with the analysis in Appendix B, the MSE of SIS is $O\left(\exp(M_{B})\right)$ with both state identification methods.
The Q-value based identification algorithm has two main advantages over the covariance testing framework. First, the algorithm is computationally inexpensive and makes it easy to include a large subset of the state space as negligible states. Second, the algorithm is easy to interpret and independent of the estimator implying it is easily used on different importance sampling frameworks such as per-decision importance sampling and incremental importance sampling.
6 State-based importance sampling variants
While so far we have focused on ordinary importance sampling, it is easy to extend SIS to use more advanced OPE estimators as its base-estimator.
First, methods such as PDIS and INCRIS are even the various weighted counterparts of these algorithms are easily extended by analogously setting the action probability ratio to one for negligible states.
Second, doubly robust off-policy evaluation [6, 21] can also benefit from state-based importance sampling with only minor adjustments. The model of the MDP and the time-specific Q-value used in the Q-value based search algorithm is based on the same formalisms that underly doubly robust (DR) off-policy evaluation. While traditional DR uses an importance weight $w_{t}^{i}=\frac{1}{n}\prod_{j=1}^{t}\frac{\pi_{e}(a_{t}^{i}|s_{t}^{i})}{\pi_{b}(a_{t}^{i}|s_{t}^{i})}$ for a trajectory $i$ at time $t$ and $n$ total trajectories, doubly robust state-based importance sampling (DRSIS) uses the importance weight $w_{t}^{i}=\frac{1}{n}\prod_{j=1}^{t}\frac{\pi_{e}^{A}(a_{t}^{i}|s_{t}^{i})}{\pi_{b}(a_{t}^{i}|s_{t}^{i})}$ where $\pi_{e}^{A}$ is the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$. To illustrate the simplicity of this change, consider that DR computes the temporary variable $\rho_{t}^{i}$ such that $\rho_{t}^{i}=\pi_{e}(a_{t}^{i}|s_{t}^{i})/\pi_{b}(a_{t}^{i}|s_{t}^{i})*\rho_{t-1}^{i}$ and the only modification for DRSIS is to set $\rho_{t}^{i}=\rho_{t-1}^{i}$ if $s_{t}^{i}\in\mathcal{S}^{A}$.
Third, techniques that rely on importance sampling over state visitation frequencies such as SDRE and MIS can also benefit from a state-based formulation. For example, the stationary density ratio estimator (Eq. 1) can be turned into a state-based formulation according to
$\hat{G}_{SSDRE}=\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{H}r_{t}^{(i)}\frac{\hat{d}^{\pi_{e}^{A}}(s_{t}^{(i)})}{\hat{d}^{\pi_{b}}(s_{t}^{(i)})}\,,$ (10)
where $\hat{d}^{\pi_{e}^{A}}(s_{t}^{(i)})$ is equal to $\hat{d}^{\pi_{e}}(s_{t}^{(i)})$ for $s_{t}^{(i)}\in\mathcal{S}^{B}$ and $\hat{d}^{\pi_{b}}(s_{t}^{(i)})$ for $s_{t}^{(i)}\in\mathcal{S}^{A}$. Such techniques are polynomial in $H$ due to replacing the product over action probability ratios by a sum of state visitation density ratios. As such the variance reduction due to reducing polynomial factors is smaller than variance reductions due to reducing the exponent, as is the case in traditional IS methods. However, as the experiments in Section 7 show, this can still yield significant MSE improvements and an excellent overall performance in environments with stationary distributions.
7 Experiments
Having introduced various state-based importance sampling techniques, we now turn to validating them experimentally. Since state-based techniques drop particular states from the computation, one would expect these to be advantageous on domains where there are some states for which the action taken has no or limited impact on the return. This intuition is captured by the notion of “lift states”.
Definition 2.
Lift states. A lift state $s\in\mathcal{S}$ is a state for which $\mathcal{T}(s,a)=\mathcal{T}(s,a^{\prime})$ and $r(s,a)=r(s,a^{\prime})$ for all pairs of actions $a,a^{\prime}\in\mathcal{A}$. The set of lift states for a domain is denoted by $\mathcal{S}^{L}$. Similarly, an approximate lift state $s\in\mathcal{S}$ is a state for which $\mathcal{T}(s,a)\approx\mathcal{T}(s,a^{\prime})$ and $r(s,a)\approx r(s,a^{\prime})$ for all pairs of actions $a,a^{\prime}\in\mathcal{A}$.
Lift states intuitively capture the covariance condition (see Condition 2) since the action probabilities, and therefore their ratios, do not affect the expected return. Analogously, for approximate lift states, action probabilities have a negligible effect on the expected return.
The experimental validation of SIS comprises 4 domains. Guided by the above intuition of lift states, the first two domains include lift states in the form of a lift transporting the agent from one side to another. As state-based importance sampling techniques are expected to be general purpose, we also assess them on two alternative domains. The third and fourth domain are common RL benchmarks, namely the inventory management domain [16] and the taxi domain variant used in [8]. Code with all estimators and experiments is available at https://github.com/bossdm/ImportanceSampling.
The experiments validate state-based importance sampling when implemented on different variants of importance sampling:
  • ordinary importance sampling (IS);
  • per-decision importance sampling (PDIS; [15]);
  • incremental importance sampling (INCRIS; [5]);
  • doubly robust OPE (DR; [21]); and
  • stationary density ratio estimation (SDRE; [8]).
The corresponding state-based implementations are denoted with the prefix S; for instance, SIS and SPDIS denote state-based IS and state-based PDIS. State-based doubly robust OPE is denoted as DRSIS to help distinguish it more clearly from SDRE. With the exception of the first domain, which has short trajectory lengths, and the SDRE technique for which this is not applicable, the experiments use additional variance reduction using weighted importance sampling [14], a technique which divides by the sum of importance weights rather than the number of trajectories. Weighted estimators are given the prefix W; for instance, WIS denotes weighted IS.
Neglible state identification parameters are fixed for each domain. Covariance testing is applied with $\epsilon_{s}=\epsilon=0.01$ in all the domains except the taxi domain for which the covariance test is too expensive to compute. The Q-value based identification is suitable for all the domains and the parameter settings are $\epsilon=1$ in the grid world lift domains, $\epsilon=50$ in the inventory management domain, and $\epsilon=2$ in the taxi domain. The effectiveness of the negligible state identification is demonstrated on the first two domains, where the performance based on state identification methods can be compared to the performance based on the lift states of these domains. The comparison for traditional importance sampling includes SIS (Lift states), which uses the lift states of the domain (i.e., where $\mathcal{S}^{A}=\mathcal{S}^{L}$); SIS (Covariance testing), which uses the negligible states identified by covariance testing; and SIS (Q-based), which uses the negligible states identified using Q-values. Similarly, we evaluate the effectiveness of state identification for the doubly robust OPE method, which are analgously denoted as SDR (Lift states), SDR (Covariance testing), and SIS (Q-based). The experiments demonstrate the Q-based identification method to correctly identify lift states and yield higher performance, which is why by default the Q-based identification is used unless indicated by the terms in parentheses.
7.1 Deterministic lift domain
The deterministic lift domain consists of consecutive lift states organised symmetrically around the starting point and with deterministic state transition dynamics. The domain involves an agent moving on a straight line, with the state space $\mathcal{S}=\{-b,\dots,-1,0,1,\dots,b\}$ defined by the bound $b$ representing the coordinates along the line and the action space $\mathcal{A}=\{\texttt{left},\texttt{right}\}$ representing left or right movements on the line. The agent starts from a central square with coordinate $s=0$ and each action moves one step left $T(s,\texttt{left},s-1)=1$ or one step right $T(s,\texttt{left},s+1)=1$ each with probability $1$ except from lift states. The episode stops when the agent hits the bound, then the terminal reward of $r=-c$ is given for if $s=-c$ and $r=c$ if $s=c$. All non-terminal rewards are $r=-1$ to penalise the length of the path. In this domain, lift states are introduced in the states $\mathcal{S}^{L}=\{-c+2,\dots,-1,1,\dots,c-2\}$, with $s\in\mathcal{S}^{L}\cap\mathbb{N}^{-}$ deterministically moving the agent further left (i.e. $T(s,\texttt{left},s-1)=T(s,\texttt{right},s-1)=1$) and $s\in\mathcal{S}^{L}\cap\mathbb{N}^{+}$ automatically moving the agent further right (i.e. $T(s,\texttt{left},s+1)=T(s,\texttt{right},s+1)=1$). The experiments manipulate this bound $c$ in $\{3,4,5,6,7,8\}$, which leads to results for domain sizes of $|\mathcal{S}|\in\{7,9,11,13,15,17\}$ and corresponds to an increasing effective horizon.
Analysing the different negligible state set identification methods reveals that SIS (Q-based) identifies the lift states as negligible states and therefore yields the same performance as SIS (Lift states). Table 5 in Appendix C demonstrates the superior performance of SIS (Q-based) and SIS (Lift states) compared to IS, which never drops any states, and SIS (Covariance testing), which typically drops two lift states (the maximal cardinality given the computational budget).
An analysis comparing state-based estimators to their traditional counterparts (see Table 1) demonstrates a near universal benefit in the deterministic domain, with the only exception being SDRE. SIS is the best performing estimator across all domain sizes, followed by IS and SPDIS depending on the domain size. SDRE and SSDRE have particularly poor estimates because the domain does not have a stationary distribution. Analysis of the residuals (see Figure 1a) also shows that: a) state-based methods have successfully reduced the variance compared to their traditional counterparts; b) SPDIS strongly reduces the residual of PDIS for large domain sizes; and c) INCRIS and SINCRIS have large residuals, which implies that dropping time steps may not be suitable for the deterministic domain (e.g. the first time step always determines the final reward).
Due to the deterministic domain being easy to model, the doubly robust methods have MSEs close to zero in this domain so are not further analysed.
(a) Deterministic lift domain
(b) Stochastic lift domain
Figure 1: Residuals ($y$-axis), defined as $\hat{G}-\mathcal{G}$, as a function of the domain size ($x$-axis) in the lift domains. Estimates are based on 1,000 episodes. Residuals are represented by their mean $\pm$ standard error over 50 independent runs for the deterministic lift domain and over 200 independent runs for the stochastic domain.
Table 1: Mean squared error (MSE) for off-policy evaluation estimators of the expected return in the deterministic lift domain based on 50 independent runs for each domain size. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size. All SIS estimators shown use the Q-value based algorithm for negligible state set identification.
(a) 100 episodes
Domain size
IS
SIS
PDIS
SPDIS
INCRIS
SINCRIS
SDRE
SSDRE
7 0.0477 0.0324
0.3732
0.2393
3.9565
3.9489
309.0670
309.9788
9 0.1084 0.0285
1.3757
0.3131
6.3628
6.2632
300.3519
304.1609
11 0.3165 0.0338
6.1956
0.6621
9.0336
8.4515
461.0101
474.6804
13 0.5054 0.0246
14.1812
0.6062
17.2282
12.4712
412.2747
430.6147
15
0.8874
0.0273
32.1658
0.8837
29.8467
15.3794
553.0677
582.1068
17
2.0117
0.0250
99.5122
0.8980
39.8329
17.1526
558.1893
587.1945
(b) 1,000 episodes
Domain size
IS
SIS
PDIS
SPDIS
INCRIS
SINCRIS
SDRE
SSDRE
7 0.0067 0.0019
0.0398
0.0121
4.0024
4.0007
532.7567
536.3787
9 0.0153 0.0019
0.1766
0.0215
6.1964
6.1765
577.3089
584.4657
11 0.0353 0.0029
0.7087
0.0600
8.9408
8.9225
674.0772
680.1959
13
0.0691
0.0026
1.8038
0.0688
12.5475
12.4779
652.6004
660.2387
15
0.1389
0.0033
5.3474
0.1378
16.1902
16.1558
803.3450
813.4997
17
0.3026
0.0031
15.1566
0.1303
23.7084
20.1534
758.2237
764.6283
7.2 Stochastic lift domain
In addition to the deterministic lift domain, we also show a stochastic lift domain in which the evaluation policy and the state transitions are stochastic rather than deterministic. The evaluation policy $\pi_{e}$ takes the best action with probability $1-\delta$ and the other action with probability $\delta$, and the state transitions dynamics are formulated as $\mathcal{T}^{\prime}(s,a,s^{\prime})=1-\delta$ where $\mathcal{T}(s,a,s^{\prime})=1$ and $\mathcal{T}^{\prime}(s,a,s^{\prime})=\delta$ otherwise – this includes lift states and non-lift states. The stochastic dynamics imply that the agent does not necessarily proceed forwards through the lift but may at some times go from one side of the line to the other side of the line. The true expected return is now negative due to typically not being able to take a straight path towards the goal, and the larger the domain the stronger this negative effect becomes.
Table 2: Mean squared error (MSE) for off-policy evaluation estimators of the expected return in the stochastic lift domain based on 200 independent runs for each domain size. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each domain size.
(a) 100 episodes
Domain size
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
7
11.8501
9.9437
7.9853
6.8646
1.6110 0.9388
141.2766
143.0795
6.1663
5.6135
9
10.4314
8.8458
6.3404
5.3060
0.5713 0.6468
163.2390
168.3226
4.0534
3.9546
11
11.5886
7.6150
6.5636
4.6576
0.8474 1.4792
161.3365
163.2882
3.3082
3.1341
13
10.3529
11.1001
7.8276
5.2161
2.3283
3.1879
190.9451
197.5445
2.9028 2.6901
15
10.1513
8.2299
8.6465
3.9065
6.9703
6.5948
213.4919
222.7848
2.3015 1.9605
17
10.5048
5.2746
12.9510
2.9489
12.5488
10.0513
246.1740
254.5904
1.9964 1.4650
(b) 1,000 episodes
Domain size
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
7
10.2238
8.4832
6.8271
5.3853
1.0958 0.2618
208.7413
209.8354
5.2572
5.0844
9
9.4737
7.7326
5.3924
4.0441
0.2973 0.6929
243.1366
245.1004
2.8067
2.7486
11
9.7285
6.2044
5.7219
3.0552
0.2002
2.3050
250.3149
251.5843
2.3016 1.9838
13
8.4436
9.4967
4.4449
3.2916
0.8958
4.8960
266.7911
268.6634
1.3459 1.3354
15
18.0453
4.2831
6.2259
1.7004
2.6550
9.0162
280.6434
282.6849
1.2258 0.8797
17
8.2474
4.8129
5.2761
1.8151
4.7157
13.0056
305.7465
307.9191
0.8385 0.4820
Results for state set identification are analogous to the deterministic lift domain: WSIS (Q-based) is successful at identifying the lift states as negligible states and together with WSIS (Lift states) has a superior performance compared to WSIS (Covariance testing) and WIS (see Table 6 in Appendix C); and WSIS (Covariance testing) typically drops two lift states (the maximal cardinality given the computational budget). Table 7 demonstrates how dropping states is also beneficial for doubly robust methods and again the Q-based state identification works to identify the lift states.
An analysis comparing state-based estimators to their traditional counterparts (see Table 2) demonstrates a benefit in the stochastic domain. This time doubly robust methods are also analysed as the error of the DR estimator is now non-zero. WSIS and WSPDIS consistently outperform their traditional counterparts. For large domain sizes, WDRSIS is the best performing estimator followed by WDR and WSPDIS. It is notable that state-based methods are not adversely affected by the domain size, which is explained by all the lift states being dropped and therefore the maximal number of steps $M_{B}$ in non-negligible states remaining constant regardless of domain size. For small domain sizes, WINCRIS and WSINCRIS have the best performance, indicating the benefit of dropping time steps; this is a clear difference compared to the deterministic domain, which is because now a) trajectories are much longer, requiring more variance reduction; and b) both the length of the trajectory and the ability of the agent to switch between the different lifts reduces the importance of any given time step, making time steps more amenable to being dropped. The lower performance of WINCRIS and WSINCRIS in larger domain sizes may be attributed to the maximal number of time steps; that is, it is not computationally feasible to check all $k\in[1,H)$ so a setting of $k\in[1,10]$ was used instead. Figure 1b further demonstrates that the state-based methods have successfully reduced the variance compared to their traditional counterparts except for WINCRIS, which has already significantly reduced the variance. As in the deterministic lift domain, the poor performance of SDRE estimators in the stochastic lift domain is attributed to the domain not having a defined stationary state distribution.
7.3 Inventory management
The third domain investigated is the inventory management problem [16]. The task of the agent in the inventory management problem is to purchase items to make optimal profits selling the items. In doing so, the agent must balance supply with demand while having at most $S$ items in the inventory. The state is the current inventory while the action is the added chosen number of items to supply. States corresponding to an almost full inventory (i.e. close to $S$ items) are not pure lift states but similar to lift states the actions chosen are unimportant – that is, the inventory will be close to or equal to $S$ regardless of the action. For each item, the purchasing cost is set to $2.49$, the sale price to $3.99$, and the holding cost to $0.03$. Demand for items follows a Gaussian with mean $S/4$ and standard deviation $S/6$, and $S=10$ in the experiments. Each episode consists of 100 steps, and consistent with the rest of our experiments an undiscounted setting ($\gamma=1$) is considered. The evaluation policy, $\pi_{e}$, was trained for 500 episodes on the MDP and has an expected return of $\mathcal{G}=1471.1041$. The behaviour policy, $\pi_{b}$, is uniform random. As the domain is high variance due to the varying demand, we opt for a relatively high number of 50 experiment runs and instead of manipulating the effective horizon we analyse the improvement in performance across different sample sizes (i.e. numbers of episodes).
Table 3: Normalised mean squared error (MSE) for Monte Carlo estimators of the expected return in the inventory management domain based on 50 independent runs for each number of episodes. Normalisation is based on division by the squared range. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each number of episodes.
Episodes
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
100
0.0490
0.0088
0.0278
0.0074
0.0641
0.0109
0.0160
0.0077
0.0173
0.0138
250
0.0651
0.0080
0.0335
0.0080
0.0546
0.0118
0.0161
0.0070
0.0132
0.0137
500
0.0652
0.0113
0.0340
0.0105
0.0456
0.0124
0.0163
0.0067 0.0098
0.0114
1000
0.0820
0.0122
0.0293
0.0119
0.0313
0.0128
0.0162
0.0066 0.0117 0.0104
An analysis comparing state-based estimators to their traditional counterparts (see Table 3) demonstrates a near universal benefit in the inventory management domain, with 2– to 8–fold improvements being observed (see Table 3), with the exception of DR which already has high accuracy. WSPDIS performs best for 100 episodes while SSDRE performs best for 250 to 1,000 episodes. The high performance of SSDRE is attributed to a clearly defined stationary distribution. The performance of SSDRE and SDRSIS depending on the sample size suggests that these state-based methods provide improved convergence and efficiency. Other high-performing algorithms are WSIS, WSINCRIS, WDR, and WDRSIS. Figure 2a demonstrates that state-based techniques reduce the variance and yield a smaller residual.
Figure 2: Residuals ($y$-axis), defined as $\hat{G}-\mathcal{G}$, as a function of the number of episodes ($x$-axis) in the inventory management domain. Residuals are represented by their mean $\pm$ standard error over 50 independent runs.
7.4 Taxi
The fourth domain is the variant of the taxi domain mentioned in Liu et al. (2018) [8], which has a much larger state space with $|\mathcal{S}|=2000$ and can have infinite horizon. The domain includes a taxi agent that has to pick up and drop off passengers who arrive at random corners in a $5\times 5$ grid based on $6$ actions, namely going a step north, west, south, east, picking up or dropping off a passenger. The taxi agent receives a reward of $20$ when picking up a passenger or dropping off a passenger at the correct location, and a reward of $-1$ otherwise. The implementation of the domain based on code of Liu et al. (2018) [8] (see {https://github.com/zt95/infinite-horizon-off-policy-estimation/tree/master/taxi}). The evaluation policy represents a policy formed after 1,000 iterations of Q-learning and the behaviour policy represents a policy formed after 950 iterations of Q-learning. The domain is of interest to investigate the scalability to large state-action spaces and large effective horizon, and it is of interest to manipulate the effective horizon in an additional domain which in contrast to the lift domains has a stationary distribution. Due to the increased computational expense, the number of runs and the number of episodes per run are limited to 20 and 100, respectively.
As in the previous domains, the results for the taxi domain (see Table 4) also show a near universal benefit for state-based estimators compared to the traditional counterparts, although the effect is not as pronounced as for example in the inventory management domain. Analysis of the effective horizon, i.e. the episode length, shows that for an effective horizon of 10, WSIS is the highest performing followed by WSINCRIS and WINCRIS while for effective horizons between 50 and 1,000, the SSDRE is the highest performer followed by SDRE and WSPDIS. The high performance of SDRE based methods, and their relative improvement across effective horizon, is attributed to their MSE being polynomial in the effective horizon whereas the MSE of traditional importance sampling techniques based on products of action probability ratios is exponential in the effective horizon. The stronger dependency of IS, PDIS, and INCRIS on the effective horizon is also demonstrated in Figure 3.
Figure 3: Normalised residuals ($y$-axis), defined as $(\hat{G}-\mathcal{G})/H$, as a function of the effective horizon ($H$; $x$-axis) of the taxi domain. Residuals are represented by their mean $\pm$ standard error over 20 independent runs.
Table 4: Mean squared error (MSE) for Monte Carlo estimators of the expected return in the inventory management domain based on 50 independent runs for each effective horizon. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each effective horizon.
Effective horizon
WIS
WSIS
WPDIS
WSPDIS
WINCRIS
WSINCRIS
SDRE
SSDRE
WDR
WDRSIS
10
0.6080
0.4525
0.7234
0.5084
0.4896 0.4767
1.2764
1.2790
53.2003
53.1011
50
15.8227
15.9281
11.6239
10.7777
35.6350
35.8282
9.4938 9.4435
188.8620
187.3876
250
440.6435
473.4986
281.9518
273.3647
1753.7938
1752.6067
25.3142 25.2920
993.3100
1055.8603
1000
8233.7773
7948.8306
4876.1759
4652.7797
27675.6630
27635.4627
231.6101 231.4313
11982.1668
12176.5372
8 Conclusion
Off-policy reinforcement learning comes with the significant challenge of high-variance estimates for the expected return. This paper introduces state-based importance sampling, a class of off-policy evaluation techniques which reduces the variance of importance sampling by decreasing the effective trajectory length through the elimination of “negligible states” for which the chosen action has no impact on the future expected return. Theoretical results on ordinary importance sampling demonstrate how the exponential dependence of the variance and MSE on $H$ can be improved into an exponential dependence on the maximal sub-trajectory length of non-negligible states. We show that either a covariance testing or a Q-value based identification algorithm can be used to identify such negligible states and that IS can be integrated with a variety of off-policy evaluation methods, including per-decision importance sampling, incremental importance sampling, weighted importance sampling, stationary density ratio estimation, and doubly robust estimation. Due to the various base-estimators’ differing assumptions, state-based techniques can be applied to a wide variety of domains by selecting a suitable base-estimator. Experimental results demonstrate the benefits of state-based methods as compared to their traditional counterparts on four distinct domains. Future work directions include theoretical studies on the presence of negligible states, implementing policy optimisation techniques based on state-based importance sampling, and exploring the links with other off-policy evaluation techniques.
Acknowledgements
David M. Bossens was supported by the UKRI Trustworthy Autonomous Systems Hub, EP/V00784X/1. Philip S. Thomas was supported by the National Science Foundation under grant no. CCF-2018372. We also thank Shreyas Chaudhari for comments on a draft of the paper.
Appendix A Appendix A: Proof of $\mathbb{E}[A]=1$
Denoting $P_{1}$ as the starting distribution, $\tau:=s_{1},a_{1}\dots,s_{H},a_{H}$, $\tau_{s}:=s_{1},\dots,s_{H}$, and $\tau_{a}:=a_{1},\dots,a_{H}$ and $\pi_{e}^{A}$ as the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$, we have
[…]
$=\int_{\tau}\pi_{e}^{A}(\tau_{a}|\tau_{s})/\pi_{b}(\tau_{a}|\tau_{s})$
$\hskip 28.45274ptP_{1}(s_{1})\prod_{t=1}^{H}\pi_{b}(a_{t}|s_{t})\mathcal{T}(s_{t},a_{t},s_{t+1})$
$=\int_{\tau}\pi_{e}^{A}(\tau_{a}|\tau_{s})P_{1}(s_{1})\prod_{t=1}^{H}\mathcal{T}(s_{t},a_{t},s_{t+1})d\tau$
[…]
$=1\,.$
Appendix B: Analysis of MSE under inaccurate state-set identification
The variance reduction of state-based methods depends crucially on the state-set identification. In covariance testing, the covariance estimate may be highly inaccurate, and in Q-value based identification, the model on which the Q-value is based may be highly inaccurate. Below we compare the true MSE of the IS estimator to that of the covariance testing based SIS estimator under the assumption that the state-identification algorithm has identified a state-set that is not negligible at all but instead has an arbitrary covariance with the return estimate (thereby yielding a significant bias in the estimate).
Note that the true MSE of ordinary importance sampling is upper bounded by
$\text{MSE}(\hat{G}_{IS})$ $\leq\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{H}\right)^{2}$
$=O(\exp(H))\,.$
While for SIS the chosen state-set has been observed to be $\epsilon$-negligible and the covariance and variance have been estimated, for the sake of the argument assume that the population covariance and variance may take on any value. Via Cauchy-Schwartz inequality and Popoviciu’s inequality with $G\in[0,Hr_{\text{max}}]$ and $A,B\in[0,\rho_{\text{max}}]$, we have
$|\text{Cov}(A,BG)|\leq\text{Var}(A)\text{Var}(BG)\leq\frac{\rho_{\text{max}}^{4}H^{2}r_{\text{max}}^{2}}{16}\,.$
The MSE of SIS is therefore given by
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))$ $=\text{Var}(\hat{G}_{SIS}(\mathcal{S}^{A}))+\text{Cov}(A,BG)^{2}$
$\leq\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}+\frac{\rho_{\text{max}}^{8}H^{4}r_{\text{max}}^{4}}{256}$
$=O(\exp(M_{B}))\,.$
Therefore the SIS has a significantly improved upper bound on the true MSE provided the maximal number of steps in $S^{B}$ is significantly smaller than $H$. If, however, the state-set identification was accurate, such that $\text{Cov}(A,BG)<\epsilon$, then any reduction in variance implies via the bias-variance decomposition that SIS has successfully optimised the true MSE compared to IS.
Appendix C: Proof of limited bias of Q-based state identification
Let $s_{t}\in S^{A}=\{s\in\mathcal{S}:|\hat{Q}_{t^{\prime}}(s,a)-\hat{Q}_{t^{\prime}}(s,a^{\prime})|<\epsilon\forall a,a^{\prime}\in\mathcal{A}\forall t^{\prime}\in[H]\}$. This implies that the SIS estimator has the following expected return from starting state $s_{t}$
$\mathbb{E}_{\pi_{b}}[\hat{G}_{SIS}|s_{t}]$ $=\mathbb{E}_{\pi_{b}}[\rho_{t+1:t+H}G_{t}|s_{t}]$
$=\sum_{a}\pi_{b}(a|s_{t})\left(r(s_{t},a)+\mathbb{E}_{\pi_{b}}[\rho_{t+1:t+H}G_{t+1}|s_{t}]\right)$
$=\sum_{a}\pi_{b}(a|s_{t})\left(r(s_{t},a)+Q_{t+1}(s_{t},a)\right)$
$=\sum_{a}\pi_{b}(a|s_{t})Q_{t}(s_{t},a)$
$\approx\sum_{a}\pi_{e}(a|s_{t})Q_{t}(s_{t},a)$
$=V_{t}(s_{t})$
$=\mathbb{E}_{\pi_{e}}[G_{t}|s_{t}]\,,$
where the approximate equality follows since $|\hat{Q}_{t}(s,a)-\hat{Q}_{t}(s,a^{\prime})|<\epsilon$ implies $Q_{t}(s,a)\approx Q_{t}(s,a^{\prime})$ for all $a,a^{\prime}\in\mathcal{A}$. Under a perfect model, the bias of the estimator comes only from the approximate equality, in which case the absolute bias induced by state being dropped is at most $\epsilon$. Under an erroneous model, which yields $|\hat{Q}_{t}(s,a)-Q_{t}(s,a)|<\epsilon_{2}$ for all $t$ and all $a\in\mathcal{A}$, the absolute bias induced by a state being dropped is at most $\epsilon+\epsilon_{2}$. Repeatedly applying this argument for a maximal number of $M_{A}$ dropped weights in a trajectory, the absolute bias induced is at most $M_{A}(\epsilon+\epsilon_{2})$. Filling in $O\left(\text{poly}(H)\right)$ as the bias and $\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}$ as the variance in the bias-variance decomposition, the MSE is bounded by $O\left(\exp(M_{B})\right)$. ∎
Appendix D: Comparison of state set identification algorithms
Table 5: Mean squared error (MSE) comparing state-based importance sampling methods with different negligible state sets on 50 independent runs of the deterministic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each number of episodes; in case of a tie only the best estimators are highlighted in bold underlined.
IS
SIS
(Lift states)
SIS
(Covariance testing)
SIS
(Q-based)
Domain size
7
0.0067
0.0019
0.0056
0.0019
9
0.0153
0.0019
0.0138
0.0019
11
0.0353
0.0029
0.0345
0.0029
13
0.0691
0.0026
0.0705
0.0026
15
0.1389
0.0033
0.1386
0.0033
17
0.3026
0.0031
0.3035
0.0031
Table 6: Mean squared error (MSE) comparing state-based importance sampling methods with different negligible state sets on 200 independent runs of the stochastic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size; in case of a tie only the best estimators are highlighted in bold underlined.
WIS
WSIS
(Lift states)
WSIS
(Covariance testing)
WSIS
(Q-based)
Domain size
7
10.2238
8.4832
10.2186
8.4832
9
9.4737
7.7326
9.4489
7.7326
11
9.7285
6.2044
9.6775
6.2044
13 8.4436
9.5003
8.4058
9.4967
15
18.0453
4.2803
17.9633
4.2831
17
8.2474
4.7482
7.7646
4.8129
Table 7: Mean squared error (MSE) comparing doubly robust methods with different negligible state sets on 200 independent runs of the stochastic lift domain for each domain size. Estimates are obtained based on 1,000 episodes. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size; in case of a tie only the best estimators are highlighted in bold underlined.
WDR
WDRSIS
(Lift states)
WDRSIS
(Covariance testing)
WDRSIS
(Q-based)
Domain size
7
5.2572
5.0844
5.2662
5.0844
9
2.8067
2.7486
2.8235
2.7486
11
2.3016
1.9838
2.3058
1.9838
13
1.3459
1.3359
1.3612
1.3354
15
1.2258
0.8788
1.1675
0.8797
17
0.8385
0.4817
0.6620
0.4820